probability of hitting a target by Ais7/10and by B is 4/10 both are independent events.find probability of hitting the target by B ,target is already hit?

Grade:12

2 Answers

Arun
4 years ago
Dear student 
 
target hitting probability =  7/10 * 6/10 + 3/10 * 4/10 + 7/10 * 4/10 = 82/100
 
now probability oh hitting target B given target has been hit already = 40/ 100 / 82/100 = 40/82
 
= 20/41
Chandrasekhar Sarma
one year ago
Probability that A hits the target = P(A) = 7/10.
Probability that B hits the target = P(B) = 4/10
Probability that target is hit = P(A+B) = P(A) + P(B)- P(AB)=P(A)+P(B)-P(A).P(B)=7/10+4/10-7/10*4/10=82/100
P((B/(A+B)) =P(B.(A+B))/P(A+B)= P(B)/P(A+B)=(4/10)/(82/100)=20/41.